OpenStudy (anonymous):

The perimeter of a rectangle is 15 cm, and its dimensions are tripled. What is the perimeter of the larger rectangle?

OpenStudy (helder_edwin):



OpenStudy (helder_edwin):

what is the perimeter of this rectangle?

OpenStudy (campbell_st):

the ratio of lengths is 1:3 so \[\frac{1}{3} = \frac{15}{?}.....? = \frac{3 \times15}{1}\]

OpenStudy (anonymous):

is it 45 ???

OpenStudy (helder_edwin):

yes